# HandSSCA: 3D Hand Mesh Reconstruction with State Space Channel Attention from RGB images

### v2

@@ Abstract
Reconstructing a hand mesh from a single RGB image is a challenging task because hands are often occluded by objects. Most previous works attempted to introduce more additional information and adopt attention mechanisms to improve 3D reconstruction results, but it would increased computational complexity. This observation prompts us to propose a new and concise architecture while improving computational efficiency. In this work, we propose a simple and effective 3D hand mesh reconstruction network HandSSCA, which is the first to incorporate state space modeling into the field of hand pose estimation. In the network, we have designed a novel state space channel attention module that extends the effective sensory field, extracts hand features in the spatial dimension, and enhances hand regional features in the channel dimension. This design helps to reconstruct a complete and detailed hand mesh. Extensive experiments conducted on well-known datasets featuring challenging hand-object occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandSSCA achieves state-of-the-art performance while maintaining a minimal parameter count.
@@ 1 Introduction
Reconstruction of 3D hand pose from a single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction. In order to reduce the cost of use in above areas, 3D hand pose estimation based on RGB images has attracted much attention due to its low cost and computational friendliness, which provides new solutions for practical applications. Influenced by the rapid development of deep learning, many excellent works Hasson2019LearningJR; Boukhayma20193DHS; Choutas2020MonocularEB; Hampali2021KeypointTS; Wang_2023; park2022handoccnet; li2022interacting have emerged in this field. These works focus on issues related to hand-object occlusion and the loss of local information. Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when there is a lack of prior knowledge.
@@
The previous methods can be broadly categorized into two groups, one is by incorporating additional a priori knowledge. For example, Zimm et al. zimmermann2017learning obtained hand region features via a hand segmentation network. And Wang et al. Wang_2023 added keypoint coordinates as strong a priori information to constrain hand pose. These methods can enhance pose estimation performance, but they lack essential a priori information in practical applications. Another category is to learn the intrinsic connection between image features through deep networks, Park et al. park2022handoccnet extracts the complete hand information by distinguishing the features. jiao et al. https://doi.org/10.1049/ell2.13040 uses the Transformer architecture to obtain the global information and expects to learn the hand feature mapping from the occluded part. These methods tend to mine the association between image features, but there are still some limitations. First, the Attention mechanism leads to excessive computational complexity. Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure.1(b).
@@
Based on the above observations, we propose a simple and effective 3D hand reconstruction network HandSSCA. This framework for the first time incorporates state space model liu2024vmamba; ma2024umamba; ruan2024vmune into the field of hand pose estimation. At the same time, combining the idea of spatial channel feature fusion hu2019squeezeandexcitation, State Space Channel Attention (SSCA) was constructed, which emphasizes the contribution of channel dimension features to image features, as shown in Fig.3. This not only increases the range of effective receptive fields in the network, but also effectively avoids the computational complexity of traditional attention square steps.Specifically, our main contributions are as follows:
@@
(1) We propose a hand pose estimation model called HandSSCA, which introduces the concept of state space model to the field of hand pose estimation for the first time. This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
@@
(2) We propose a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning. And the state space channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
@@
Figure 1: Effective receptive fields (from left to right, FPN, Self-Attention, State Space Attention(SSA) and State Space Channel Attention(SSCA), respectively).
@@ 2.1 State Space Model
With the proposal of attention mechanism vaswani2023attention, it makes more models can effectively focus on global information mapping, but its quadratic computational complexity leads to computational inefficiency in the case of long sequences. In the field of natural language processing, Albert Gu et al. gu2022efficiently proposed the S4 model to optimize the state space model (SSM) by simulation in order to solve the problems of computational efficiency and long sequence dependency. However, the S4 model would process all inputs in exactly the same way, which is not in line with the practical intuition that each sequence should receive different attention. Therefore, Albert Gu et al. gu2023mamba introduced a selection mechanism to propose the Mamba architecture (S6) to learn the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences. Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to linear. Influenced by this idea, Yue Liu et al. ma2024umamba introduced it to vision by combining the VIT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity. This approach achieves linear complexity without sacrificing the global receptive field, as shown in Figure.1(c). In addition, some researchers ruan2024vmunet; hu2019squeezeandexcitation have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information.
2.2 Additional information based Methods
It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations. In recent years, some scholars Wang_2023; jiang2023a2jtransformer combined human a priori knowledge to guide the network to focus on the target region, Wang et al. Wang_2023 proposed HandGCAT, which learns the a priori knowledge of the hand joint positions through the Knowledge-Guided Graph Convolution module, and then through the Cross-Attention Transformer to guide the network to inject joint information into the occluded region. In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose, therefore, Qi et al. qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
2.3 Attention based Methods
Attentional mechanisms have achieved great success in the field of deep learning, thanks to their ability to map long-term contextual relevance, and in the task of hand pose estimation, occluded hand features can be learned through global mapping, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing. In establishing the global context of hand-object interaction, Pavlakos et al. pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the VIT architecture.Cho et al. cho2022crossattention obtain the features related to the region of interest from the occluded portion through cross-attention. Park et al. park2022handoccnet make full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
@@ 2.3 Attention based Methods
Figure 2: Overall architecture of HandSSCA including Trunk, SSCA and Regressor.Our HandSSCA uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer. It then uses SSCA to obtain hand region features by pairwise spatial and channel scanning. Finally, the regressor obtains keypoint information by heatmap encoding and outputs MANO parameters. The final 3D hand mesh is obtained by forwarding the MANO parameters to the MANO layer.
3 Method
@@ 3.1 Overview
The overall network architecture is shown in Figure.2 and consists of Backbone, SSCA and Regressor components. Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$. First, the image information is extracted by ResNet-50 and the multi-scale features are fused layer by layer to get the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$. Then, it is fed into the SSCA module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain hand features $H\in\mathbb{R}^{256\times 32\times 32}$. Finally, the hand features $H$ are passed to Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates and output the MANO parameters through the FC layer to obtain the final 3D hand mesh.
@@ 3.2 Backbone
In the feature extraction part we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature. Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$. In the subsequent layers, each layer of image features $F_{i}\in\frac{H_{i}}{2^{i}}\times\frac{W_{i}}{2^{i}}\times C_{i},i\in(1,2,3,4),C_{i}\in(256,512,1024,2048)$ is obtained by downsampling. Then, up-sampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of features of size $64\times 64\times 256$. Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the image information through average pooling.
3.3 State Space Channel Attention
@@
S4ND nguyen2022s4nd was the first study to introduce state space models into the field of vision by enabling it to receive 2D image information through a simple extension of the S4 model. The S6 module proposed by Gu et al. gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selectable scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity. The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{C}^{N\times N},B,C\in\mathbb{C}^{N},D\in\mathbb{C}^{1}$ with the following equation:
$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\\ \mathrm{y(t)=Ch(t)+Dx(t)}$ (1)
In order to achieve computationally efficient operations zimmermann2019freihand, the above ordinary differential equations need to be transformed into discrete functions. Specifically, when the input size is $X_{i}\in\mathbb{R}^{\mathrm{L\times D}}$, the discretization is performed following the following rule:
@@
where $\mathrm{B,C\in\mathbb{C}^{D\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{D}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series ref16:
@@
Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B\in\mathbb{R}^{B\times L\times N}\quad,C\in\mathbb{R}^{B\times L\times N}$, and $\Delta\in\mathbb{R}^{B\times L\times D}$ can be obtained from inputs of size $X\in\mathbb{R}^{B\times L\times D}$. this means that S6 can sense contextual information from the inputs and update the weights.
Space select.
@@
Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module. The 1-D convolution is extended to 2-D convolution in S4ND gu2021combining, and this approach leads to the loss of the ability to dynamically update weights based on context. In VMamba liu2024vmamba, rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module. This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane. Then, the sequences from each of the four directions are passed through the S6 block to compute the feature map. Finally, the four sequences are merged to obtain a new sequence of the same size as the input.
Channel select.
@@
Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al liu2024vmamba. However, in deep image features, channels contain richer and diverse contextual information, which is mostly ignored by these methods. Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Figure.3. We perform scan expansion and scan fusion on the $H\times C$ plane. First, the input features are expanded on the $H\times C$ plane and scanned from four directions to obtain the sequence $C_{1},C_{2},C_{3},C_{4}$. Then, they are fed into the S6 block to compute the contextual correlation in the sequence by Eq. (3), respectively. Finally, the sequences are merged and reconstructed into image feature maps of the same size.
@@
Image features are spatially selected by activating the mapping of each pixel in the spatial plane and the four directions of up, down, left and right to obtain local texture and structure information. In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field. We visualized the range of effective receptive fields for different attentions, as shown in Figure.1. As can be seen in Figure.1(d), this novel state space channel attention is able to maintain a larger and deeper receptive field range and global feature mapping capability with a reduced number of parameters and linear complexity. This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
Figure 3: SSCA module framework, we send the image features to parallel spatial and channel branches for scanning, and then process the spatial and channel resultant features and merge them through S6 block respectively, and finally sum them to get the 2D feature map.

### v3

# HandS3C: 3D Hand Mesh Reconstruction with State Space Spatial Channel Attention from RGB images

@@ Abstract
Reconstructing the hand mesh from one single RGB image is a challenging task because hands are often occluded by other objects. Most previous works attempt to explore more additional information and adopt attention mechanisms for improving 3D reconstruction performance, while it would increase computational complexity simultaneously. To achieve a performance-reserving architecture with high computational efficiency, in this work, we propose a simple but effective 3D hand mesh reconstruction network (i.e., HandS3C), which is the first time to incorporate state space model into the task of hand mesh reconstruction. In the network, we design a novel state-space spatial-channel attention module that extends the effective receptive field, extracts hand features in the spatial dimension, and enhances regional features of hands in the channel dimension. This helps to reconstruct a complete and detailed hand mesh. Extensive experiments conducted on well-known datasets facing heavy occlusions (such as FREIHAND, DEXYCB, and HO3D) demonstrate that our proposed HandS3C achieves state-of-the-art performance while maintaining a minimal parameters.
@@ 1 Introduction
The reconstruction of 3D hand mesh from one single RGB image has a wide range of applications in many fields, such as VR/AR, robotics, and human-computer interaction. 3D hand mesh reconstruction based on RGB images for the reconstruction has attracted much attention due to its low cost and computational friendliness. Influenced by the rapid development of deep learning, many excellent works Hasson2019LearningJR; Boukhayma20193DHS; Choutas2020MonocularEB; Hampali2021KeypointTS; Wang_2023; park2022handoccnet; li2022interacting have emerged in this field. These works focus on important issues related to hand-object occlusion and the loss of local information. Despite significant advancements, most methods still struggle to reconstruct hand poses in scenarios with substantial occlusion or when the prior knowledge is not sufficient.
@@
The previous methods can be broadly categorized into two categories. One is to incorporate additional prior knowledge. To cite few, Zimm et al. zimmermann2017learning obtained hand region features via a hand segmentation network. Wang et al. Wang_2023 added keypoint coordinates as strong prior information to constrain hand pose. These methods can enhance pose reconstruction performance, but lack essential prior information in practical applications. Another category is to learn the intrinsic connection between image features through deep networks, Park et al. park2022handoccnet extracted the complete hand information by distinguishing the features. Jiao et al. https://doi.org/10.1049/ell2.13040 used the Transformer architecture to obtain the global information and expect to learn the hand feature mapping from the occluded part. These methods tend to mine the association between image features, but there are still some limitations. First, the attention mechanism leads to excessive computational complexity. Second, an overload of global information can interfere with the effective receptive field luo2017understanding, as shown in Figure. 1 (b).
@@
Based on the above observations, we propose a simple but effective 3D hand mesh reconstruction network called as HandS3C. This framework incorporates the state space model liu2024vmamba; ma2024umamba; ruan2024vmunet into the field of hand mesh reconstruction for the first time. At the same time, combining the idea of spatial-channel feature fusion hu2019squeezeandexcitation, State Space Spatial Channel (S3C) attention module is constructed to emphasize the contribution of channel-dimensional features. This not only increases the range of effective receptive fields in the network, but also effectively avoids the square-step computational complexity of traditional attention. Specifically, our main contributions are summarized as follows:
@@
(1) We design a 3D hand mesh reconstruction network called as HandS3C, which introduces the state space model to the task of hand mesh reconstruction for the first time. This network can effectively improve hand reconstruction performance without the need for additional prior knowledge.
@@
(2) We propose a spatial and channel-wise parallel scanning approach to compensate for the missing channel information in planar scanning. The state-space spatial-channel attention module is constructed, which can enhance the effective receptive field range while maintaining a small number of parameters.
@@
Figure 1: Effective receptive fields (from left to right: FPN, Self-Attention, State Space Attention (SSA) and State-Space Spatial-Channel Attention (S3C)).
@@ 2 Related work
2.1 Additional Information based Methods
It is difficult to restore 3D hand gestures from RGB image features alone, and most researchers use heatmaps Boukhayma20193DHS; lv2021handtailor; iqbal2018hand, segmentation zimmermann2017learning; zhang2021hand, depth maps rezaei2022trihornnet; rezaei2023pushing; 10.1007/978-3-030-01270-0_15, and density maps zhao2023semisupervised; chen2021i2uvhandnet to incorporate additional information to help the network better learn hand features in occlusion situations. In recent years, some scholars Wang_2023; jiang2023a2jtransformer incorporate human prior knowledge to guide the network to focus on the target region, Wang et al. Wang_2023 proposed HandGCAT, which learns the prior knowledge of the hand joint positions through the knowledge-guided graph convolution module, and then through the cross-attention Transformer to guide the network to inject joint information into the occluded region. In addition, other scholars Hasson2019LearningJR; wang2022interacting; qi2024hoisdf; li2022artiboost; chen2022tracking have argued that the object pose affects the hand pose. To end this, Qi et al. qi2024hoisdf introduced Signed Distance Field (SDF) to reconstruct the occluded object and the hand, and this method is able to determine the grip state of the hand by estimating the object pose and then determining the hand’s grasping state.
2.2 Attention based Methods
Attention mechanisms have achieved great success in the field of deep learning, due to their ability to map long-term contextual relevance. In the task of hand mesh reconstruction, occluded hand features can be learned through global mapping with various attention modules, which in turn can effectively solve the occlusion problem park2022handoccnet; li2022interacting; cho2022crossattention; feng2021collaborative; pavlakos2023reconstructing. In establishing the global context of hand-object interaction, Pavlakos et al. pavlakos2023reconstructing learn the mapping relationship between hand and occlusion from large-scale data through the Vision Transformer (ViT) architecture. Cho et al. cho2022crossattention obtained the features related to the region of interest from the occluded portion through cross-attention. Park et al. park2022handoccnet made full use of the information of occluded region as an enhancement of image feature of the hand and refine it to get the complete hand features.
@@ 2.3 State Space Model
With the proposal of attention mechanism vaswani2023attention, it makes more models can effectively focus on global information mapping, but its quadratic computational complexity leads to computational inefficiency in the case of long sequences. In the field of natural language processing, Gu et al. gu2022efficiently proposed the S4 model to optimize the state space model (SSM) in order to solve the problems of computational efficiency and long sequence dependency. However, the S4 model would process all inputs in the exactly same way, which is not in line with the practical intuition that each sequence should receive different attention. Therefore, Gu et al. gu2023mamba introduced a selection mechanism to the Mamba architecture (S6) for learning the parameters in the traditional state-space model, through which it is able to focus on or ignore some sequences. Compared to traditional attention computation methods, S6 enables each element in a 1-D array (e.g., a text sequence) to interact with any previously scanned samples through a compressed hidden state, effectively reducing the quadratic complexity to a linear one. Influenced by this idea, Liu et al. liu2024vmamba introduced it to vision tasks by combining the ViT model dosovitskiy2021image to transform images into sequences, and proposed a cross-scanning module to solve the problem of image orientation sensitivity. This is made possible by the selectivity of the state space model, which allows it to enhance similar features in the hidden space. Therefore, we abstract the selective state-space model and the cross-scanning module therein into a state space attention (SSA) module, as shown in Figure. 1 (c). In addition, some researchers ruan2024vmunet; ma2024umamba; Deng2024CUMambaSS have also combined state-space modeling with traditional CNN architectures to extract multi-scale global features in U-shaped structures by state-space selection modeling, and retained the deep-level information. In this paper, we introduce state space model into the field of hand mesh reconstruction for the first time , and design a state space spatial channel attention module (S3C). The module performs parallel scanning in both spatial and channel dimensions, and uses a selective state-space model to achieve feature selection in the hidden space. This method is able to extend the effective receptive field with a small number of parameters and focuses the network attention on the hand region, thus improving the accuracy of hand mesh reconstruction.
@@
Figure 2: Overall architecture of HandS3C including Backbone, S3C and Regressor. HandS3C uses Resnet50 to extract image features and fuses multiscale features by upsampling layer by layer. It then uses S3C to obtain hand region features through parallel spatial and channel scanning. Finally, the regressor obtains keypoint information by heatmap encoding and outputs MANO parameters. The final 3D hand mesh is obtained by forwarding the MANO parameters to the MANO layer.
3 Methodology
@@ 3.1 Overview
The overall network architecture is shown in Fig. 2 and consists of Backbone, S3C and Regressor components. Given a hand image $I\in\mathbb{R}^{256\times 256\times 3}$. First, the appearance information is extracted by ResNet-50 based encoder-decoder and the multi-scale features are fused layer by layer to get the original feature map $F\in\mathbb{R}^{256\times 32\times 32}$. Then, it is fed into the S3C module and scanned in both spatial and channel dimensions to compute a 2D feature map and fused to obtain enhanced hand features $H\in\mathbb{R}^{256\times 32\times 32}$. Finally, the hand features $H$ are passed to the Regressor to calculate the keypoint heat map $HM\in\mathbb{R}^{21\times 32\times 32}$ to predict the keypoint coordinates. For the final 3D hand mesh, we used the most popular MANO parametric hand MANO:SIGGRAPHASIA:2017. It is an articulated hand network generated from differentiable functions that takes as input the parameters that control the shape $\beta\in\mathbb{R}^{10}$ and pose $\theta\in\mathbb{R}^{48}$ of the hand.
@@ 3.2 Backbone
In the original feature extraction part, we use ResNet-50 he2015deep to extract image information and fuse multi-scale features layer by layer lin2017feature in an U-shape encoder-decoder structure. Specifically, when the input image size is $256\times 256\times 3$, the size of the first layer of features after ResNet-50 becomes $128\times 128\times 128$. In the subsequent layers, each layer of image features $F_{i}\in\frac{H_{i}}{2^{i}}\times\frac{W_{i}}{2^{i}}\times C_{i},i\in(1,2,3,4),C_{i}\in(256,512,1024,2048)$ is obtained by downsampling. Then, upsampling is performed and the image features $F_{i}$ are incorporated layer by layer to get the last layer of feature extraction with size of $256\times 64\times 64$. Finally, the feature mapping map $F\in\mathbb{R}^{256\times 32\times 32}$ is obtained by fusing the feature information through average pooling.
3.3 State Space Spatial Channel Attention Module
@@
S4ND nguyen2022s4nd was the first study to introduce state space models into the field of vision by enabling it to receive 2D image information through a simple extension of the S4 model. The S6 module proposed by Gu et al. gu2023mamba was shown in VMamba liu2024vmamba to be able to capture the global feature information of 2D image features through a selective scanning mechanism enabling it to capture the global feature information of the 2D image features and maintain linear complexity. The state space model expects to map the input $x(t)\in\mathbb{R}^{L}$ to the output $y(t)\in\mathbb{R}^{L}$ through the implicit state $h(t)$ and the parameter $A\in\mathbb{R}^{N\times N},B,C\in\mathbb{R}^{N},D\in\mathbb{R}^{d}$ with the following equation:
$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),\;y(t)=Ch(t)+Dx(t)$ (1)
In order to computationally achieve efficient operations gu2021combining, the ordinary differential equations above need to be transformed into discrete functions. Specifically, when the input size is $X_{i}\in\mathbb{R}^{\mathrm{L\times d}}$, the discretization is performed following the following rule:
@@
where $\mathrm{B,C\in\mathbb{R}^{d\times N}}$ and $\Delta\in\mathbb{R}^{\mathrm{d}}$. In practice, we improve the $\bar{B}$ approximation by means of a first-order Taylor series gu2023mamba:
@@
Unlike traditional spatial state models, the S6 block introduces a selective scanning mechanism, where $B,C\in\mathbb{R}^{bs\times L\times N}$, and $\Delta\in\mathbb{R}^{bs\times L\times d}$ can be obtained from inputs of size $X\in\mathbb{R}^{bs\times L\times d}$. This means that S6 can sense contextual information from the inputs and update the weights.
Spatial-wise Selective Scanning.
@@
Unlike textual information, image information contains 2D spatial information such as local texture and global structure, and it is difficult to obtain the contextual relationship of the local sense field of an image by scanning with a simple S6 module. The 1-D convolution is extended to 2-D convolution in S4ND gu2021combining, and this approach leads to the loss of the ability to dynamically update weights based on context. In VMamba liu2024vmamba, rows and columns are unfolded and scanned along in four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right, by means of a cross-scanning module. This allows any pixel to contain contextual information from all four directions and can enhance the size of the sensory field in the $H\times W$ plane. Sequences in each of the four directions are computed in the Spatial-wise Selective Scanning to obtain $B,C\in\mathbb{R}^{bs\times N\times L_{H\times W}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{H\times W}}$, which in turn computes the feature map. Finally, the four sequences are merged to obtain a new sequence with the same size as the input sequence.
Channel-wise Selective Scanning.
@@
Although spatial features such as local texture and global structure intrinsic to 2D image features are already available through Liu et al liu2024vmamba. However, in deep features, channels contain richer and diverse contextual information, which is mostly ignored by these methods. Therefore, we design a spatial and channel-based parallel scanning approach to compensate for the missing channel information in planar scanning, as shown in Fig. 3. We perform scan expansion and scan fusion on the $ch\times H$ plane. First, the input features are expanded on the $ch\times H$ plane and scanned from four directions to obtain the sequence $X_{1},X_{2},X_{3},X_{4}$. Then, in Channel-wise Selective Scanning, $B,C\in\mathbb{R}^{bs\times N\times L_{ch\times H}}$, and $\Delta\in\mathbb{R}^{bs\times d\times L_{ch\times H}}$ is calculated by Eq. 1-3 as an indication of the correlation between the feature channels. Finally, the sequences are merged and reconstructed into feature maps with the same size for images.
@@
Image features are spatially selected by activating the mapping of each pixel in the spatial plane and the four directions of up, down, left and right to obtain local texture and structure information. In channel selection, the mapping relationship between different channels can be obtained to enhance the representation of similar channels in the implicit space in order to construct a wide range of global feature mappings and to enhance the scope of the sensory field. We visualized the range of effective receptive fields for different attentions, as shown in Fig. 1. As can be seen in Fig. 1 (d), the novel state space spatial channel attention module is able to maintain a larger and deeper receptive field and global feature mapping capability with a reduced number of parameters and linear complexity. This is consistent with our motivation to expect that enhancing the effective receptive field can better capture hand-object information associations from image features.
Figure 3: S3C module. In the S3C module, image features are fed into Spatial-wise Selective Scanning and Channel-wise Selective Scanning respectively. The correlation of spatial and channel dimensions in the hidden space is computed by selective state space modeling respectively. Eventually, they are merged to activate hand region features to obtain a feature map.